OpenStudy (anonymous):

hey, how do you express sin(wt) into 1/j2 * (e^(jwt) - e^(-jwt)) its normally given to you in a maths booklet but i don't know how to get to it thank you x

OpenStudy (ash2326):

e^(jwt)=cos wt+ j sin wt e^(-jwt)=coswt - j sin wt e^(jwt)-e^(-jwt)=2j sin wt sin wt= (e^(jwt)-e^(-jwt))/2j

OpenStudy (anonymous):

thank you for the help x